OpenStudy (anonymous):

Hannah has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $28.30 How many quarters and how many nickels are in the jar? (Points : 1) 46 quarters; 104 nickels 75 quarters; 75 nickels 104 quarters; 46 nickels 95 quarters; 55 nickels

OpenStudy (anonymous):

@JReed231

OpenStudy (anonymous):

i think 75 quarters and on 75 nickels

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Amy mixes 12 liters of 16% acid solution with a 40% acid solution which results in 32% acid solution. If x represents the number of liters of 40% solution used, which equation can be used to find the number of liters of 40% acid solution in the mixture? (Points : 1) 1.92 + 0.32x = 0.4(x + 12) 3.84 + 0.4x = 0.16(x + 12) 4.8 + 0.32x = 0.4(x + 12) 1.92 + 0.4x = 0.32(x + 12)

OpenStudy (anonymous):

thanks by the way

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

Amy mixes 12 liters of 16% acid solution with a 40% acid solution which results in 32% acid solution. If x represents the number of liters of 40% solution used, which equation can be used to find the number of liters of 40% acid solution in the mixture? (Points : 1) 1.92 + 0.32x = 0.4(x + 12) 3.84 + 0.4x = 0.16(x + 12) 4.8 + 0.32x = 0.4(x + 12) 1.92 + 0.4x = 0.32(x + 12)